### v1

# A Simple and Explainable Model for Park-and-Ride Car Park Occupancy Prediction

Abstract
In a scenario of growing usage of park-and-ride facilities, understanding and predicting car park occupancy is becoming increasingly important. This study presents a model that effectively captures the occupancy patterns of park-and-ride car parks for commuters using truncated normal distributions for vehicle arrival and departure times. The objective is to develop a predictive model with minimal parameters corresponding to commuter behaviour, enabling the estimation of parking saturation and unfulfilled demand.
The proposed model successfully identifies the regular, periodic nature of commuter parking behaviour, where vehicles arrive in the morning and depart in the afternoon. It operates using aggregate data, eliminating the need for individual tracking of arrivals and departures. The model’s predictive and now-casting capabilities are demonstrated through real-world data from car parks in the Barcelona Metropolitan Area. A simple model extension furthermore enables the prediction of when a car park will reach its occupancy limit and estimates the additional spaces required to accommodate such excess demand. Thus, beyond forecasting, the model serves as a valuable tool for evaluating interventions, such as expanding parking capacity, to optimize park-and-ride facilities.
keywords:
Car Park, Occupancy Prediction, Demand estimation, Park-and-ride, Commuter Behaviour
1 Introduction
The rapid increase in energy prices and the growing awareness of the need to reduce greenhouse gas emissions have led an increasing number of commuters to shift from private to public transport. Additionally, major urban centres such as London, Barcelona, and Vienna are progressively restricting access to city centres for private vehicles or limiting free-of-charge parking by expanding regulated parking zones. In response to these developments, park-and-ride systems have become a crucial solution, allowing commuters to leave their private vehicles at designated facilities outside city limits and complete their journeys using public transport.
Effective park-and-ride systems require large parking facilities strategically located near major transport hubs, such as train stations. Accurately estimating the required capacity of these parking facilities and expanding them when demand exceeds available space is a critical aspect of urban planning. A thorough understanding of commuter behaviour and the ability to predict parking demand are essential for making informed decisions in this context.
In this work, we use data collected from a set of eight park-and-ride facilities located in the vicinity of Barcelona, Catalonia. These car parks primarily serve commuters on working days, who typically park their vehicles in the morning before transferring to public transport for work or study in the city. In the afternoon, they return to retrieve their cars and drive home.
The observed periodic patterns in car park occupancy are highly regular, making them well-suited for analysis and prediction using simple models. Our goal is to develop models that improve the planning and management of park-and-ride infrastructure. The specific objectives of our modelling approach are as follows:
  • Use as few parameters as possible.
  • Ensure that the model parameters are interpretable and correspond to behavioural metrics.
  • Enable the model to predict when a parking lot will be full.
  • Allow the model to assess the unmet demand for free parking spaces in a car park.
To achieve these goals, we will model car arrivals and departures separately, each using a different truncated normal distribution. This approach allows us to use aggregate data without the need to monitor individual arrivals and departures. Before presenting the proposed model in Section 4, we first review relevant literature in Section 2. In Section 3, we describe the dataset used in our study. Section 5 then analyzes the model’s ability to fit the observed data and evaluates its performance in both prediction and nowcasting tasks. Finally, we conclude with a discussion of our findings in Section 6.
2 Related work
The literature about smart parking solutions is vast and includes different disciplines and types of problems as can be observed in this survey paper [1]. Here, we focus on a specific problem in this area: Modeling car park occupancy. This has been treated in the literature mostly only as a prediction problem (see [2] for a recent review on the subject). Two different types of prediction problems can be distinguished: on-street and off-street (e.g. in garages and parking lots) parking prediction. We are aiming at the latter.
Notable works for car park occupancy prediction include [3], [4], and [5]. Reference [3] compares the performance of regression trees, support vector regression and a neural network with a single hidden layer for sensor-enabled car parks from San Francisco (US) and Melbourne (Australia). The study found that the least computationally intensive algorithm of the three, the regression trees with the history of the occupancy rates and the weekday as input features performed best.
Recurrent neural networks with more hidden layers have then been employed by [4] to predict car park occupancy rates of 29 car parks in Birmingham, UK.
Furthermore, the authors of [5] also give an overview of different parking occupancy prediction methods. Their study compares methods based on linear regression, support vector machine (SVM), neural networks and auto-regressive integrated moving average (ARIMA) which are applied to data from four car parks in three Chinese cities. The effect of distinguishing between weekend and weekday patterns is also analyzed. The study concludes that SVM offers stable and accurate prediction performance for almost all types and scales of parking lots while the distinction between weekday and weekend only led to improvements for medium size parking lots and office space parking.
In the context of commuter behaviour, [6] analyzed the relationship between parking space availability and commuters’ departure time choices using survey data. This study also includes a model with a bell-shaped distribution for the perceived value of choosing a specific departure time, which produces departure distributions similar to the ones we use in our models.
Patterns with multiple daily peaks, more complex than the ones we observe here have been analyzed in [7] with the aim of predicting parking space availability in a more generic setting without using an underlying behavioural model. The same aim has been pursued in [8] which models individual parking lot occupancy as a two-state stochastic process. Similar patterns have been analysed recently in [9] where the number of parked car-sharing vehicles in specific areas in the Milan metropolitan area are predicted with a Maximum Entropy modelling framework which allows identifying extreme events such as bad weather conditions or strikes.
Here, we take a slightly more ambitious approach by aiming to understand the underlying behavioural patterns. In particular, we compare our method with simple prediction techniques based on average activity profiles, as used in [10] and [11]. Previous studies have shown that historical data can effectively predict future parking occupancy rates, even with limited data [12].
This study (as does the above-mentioned reference [5]) uses among outer algorithms linear regression as prediction algorithm. We will use this method as an additional baseline to compare the predictive power of our modelling approach.
For cases where (contrary to our case) individual arrival and departure times are available a related modelling approach based on queuing models with non-homogeneous arrival rates was undertaken in [13]. Using data from truck parking lots the authors found that a queuing model of parking dynamics and a model-based prediction method can provide real-time probabilistic forecasts of future parking occupancy.
Behavioural decision models for park-and-ride facilities utilization have been analyzed in [14], whose authors use survey data to analyze how personal attributes, travel characteristics, and user intentions influence the decisions to choose or not a Park-and-Ride trip. Shortage of parking space is found to be among the top three reasons not to use Park-and-Ride facilities. This shows the importance of not only correctly predicting the demand for parking space but also the missing amount of parking space which cannot be fulfilled by the current infrastructure. None of the above-mentioned studies has tackled this problem, which our behavioural model can address.
3 Dataset Description
We use data provided by the metropolitan transport authority  [^1] Autoritat de Transport Metropolità (ATM): https://www.atm.cat/en/atm. of Barcelona (Catalonia) which has installed sensors in different park-and-ride parking locations with the aim to better understand their usage patterns and to develop a long-term strategy that allows to improve the interurban mobility from and to Barcelona. Each of the car parks is located in the vicinity of a regional train station which allows people to park their cars and change directly there to the public transportation system, working as a bridge between both public and private transportation means [15].
Figure 1: Locations of the park-and-ride car parks analyzed, all of them in the vicinity of Barcelona (Catalonia). Two of the locations, Martorell and San Quirze, have not been used due to errors in the data collection process.
Data from a total amount of 10 parking lots have been provided. Their geographical location is depicted in Figure 1. Data is continuously gathered from physical sensors placed in the entrances and exits of the car parks and stored in occupancy counters. An entrance of a vehicle increases the occupancy of the particular parking, and a departure of a vehicle decreases that counter. No data about the precise physical location where a vehicle park is recorded. This occupancy data is sampled in 30-minute internals from every parking during a 3-month period from January 1st 2020 0:00 to April 1st 2020 0:00. Note that Catalonia entered a COVID-19-induced nationwide lockdown on March 15, 2020, and the car park usage decreased significantly. We will omit the data posterior to this date in the subsequent analysis. We furthermore have removed data from two car parks (Martorell and San Quirze) due to defective sensors and storage errors which led to irregular readings.
Figure 2 gives an example of the raw data from one of the car parks (Vilanova). The observed occupancy daily pattern is the aggregate result of arrivals and departures to the car park location. In the subsequent analysis, we have manually removed days where the occupancy profiles show errors due to sensor or storage failures. An example of such days are the days from February 7 to February 9 in Figure 2 or days with are influenced by holidays or vacation periods (e.g days 1 to 3 and 6 of January in Figure 2, both periods together with the also removed COVID-19 lockdown period are shown with a grey background.
Figure 2: Example of input data from the Vilanova car park, days with grey background have been removed from the analysis due to sensor failure (7th to 9th of February), COVID-19 induced lockdown (days after March 15th) or reduced activity due to Christmas holidays (1st to 3rd and 6th of January)
This raw data already hints at the regularity in the parking behaviour. If we aggregate this data into weekly or daily activity cycles as is done in Figure 3, we can observe the regularity of the corresponding circadian patterns. The average behaviour from Monday to Thursday is nearly identical, and activity on Fridays shows a small decay.
In what follows we aggregate the data in three groups: weekdays from Monday to Thursday, Fridays, and weekends (Saturdays and Sundays)
Figure 3: Average weekly (top) and daily activity cycles (bottom) of the Vilanova car park. The daily average cycles are very similar from Monday to Thursday.
Finally in Figure 4 we look at how close the different car parks get to their capacity limit (indicated by the red vertical line). For every day (excluding the days we have filtered out as mentioned above) we record the maximum occupancy of the car park and depict the corresponding distributions. Four of the car parks (Sant Sadurni, Sant Boi, Quatre Camins and Mollet) reach their maximal occupancy limit regularly during weekdays (orange bars) and also Fridays (blue bars). In the case of the example we used above, the Vilanova car park, this is not the case. We will use this distinction and develop a model extension for the stations where the capacity limit is reached.
Figure 4: Histogram of the observed maximal occupancies for the days between 01-01-2020 and 01-04-2020 aggregated by weekdays (orange), Fridays (blue) and weekends (green) for the eight car parks used in this study. Red lines indicate the maximal capacities of the car parks.
4 Modeling Car Park Occupancy
Given the regularity of the activity cycles we have seen in Figure 3 we now present statistical models able to describe the daily occupancy patterns of an individual parking lot. The model considers the parking lot occupancy as a combination of two independent processes: arrivals and departures. At any time, the occupancy is simply the aggregated number of arrivals minus the aggregated number of departures.
We introduce two types of models: a basic model which assumes an unlimited amount of free parking spaces, and an extended model with occupancy limit. The former is suitable for parking lots which do not fill up completely whereas the latter takes into account that a parking lot may reach its capacity limit and cars arriving after this moment will have to find parking space elsewhere.
4.1 Basic Model (Unlimited Parking Spaces)
The proposed model accounts for two event types: car arrivals and departures. For simplicity, these events are assumed to be independent. Each process is modelled as a random variable following a (truncated) normal distribution. The truncation is necessary to constrain the arrival and departure times to occur within the period corresponding to one day (which we normalize to the interval $[0,1]$). The probability density functions for arrivals $\phi_{a}$ and departures $\phi_{d}$ are thus:
$\phi_{a}(t;\mu_{a},\sigma_{a})$ $=\frac{\phi(\frac{t-\mu_{a}}{\sigma_{a}})}{\sigma_{a}\cdot\left(\Phi\left(\frac{1-\mu_{a}}{\sigma_{a}}\right)-\Phi\left(\frac{-\mu_{a}}{\sigma_{a}}\right)\right)},$ (1)
$\phi_{d}(t;\mu_{d},\sigma_{d})$ $=\frac{\phi(\frac{t-\mu_{d}}{\sigma_{d}})}{\sigma_{d}\cdot\left(\Phi\left(\frac{1-\mu_{d}}{\sigma_{d}}\right)-\Phi\left(\frac{-\mu_{d}}{\sigma_{d}}\right)\right)},$ (2)
where $\phi(\cdot)$ is the probability density function of the standard normal distribution and ${\displaystyle\Phi(\cdot)}$ its cumulative distribution function (CDF). We denote the truncated CDFs as $\Phi_{a}(t;\mu_{a},\sigma_{a})$ for the arrival and $\Phi_{d}(t;\mu_{d},\sigma_{d})$ for the departure process, which are determined by parameters $(\mu_{a},\sigma_{a})$ and $(\mu_{d},\sigma_{d})$ respectively.
At any time, the occupancy at the parking lot is given by the number of arrivals minus the number of departures. We can generate a daily occupancy realization by drawing $M$ sample times from both processes $\{A_{n}\}_{n=1}^{M}\sim\phi_{a}$ and $\{D_{n}\}_{n=1}^{M}\sim\phi_{d}$ and subtracting the corresponding counting processes:
$o(t)=\sum_{n\geq 1}\mathbbm{1}_{\{t\geq A_{n}\}}-\sum_{n\geq 1}\mathbbm{1}_{\{t\geq D_{n}\}},\qquad t=[0,1],$ (3)
where $\mathbbm{1}_{\{\cdot\}}$ is the indicator function that takes value 1 when the condition is true and 0 otherwise. To know the occupancy of a car park at a given time $t$ we calculate the difference between all arrivals and departures happened before.
A (normalized) continuous approximation to Eq. (3) takes the difference between the CDFs of the arrivals $\Phi_{a}(\cdot)$ and departures $\Phi_{d}(\cdot)$:
$f_{\text{TN}}(t;\theta)$ $=\Phi_{a}(t;\mu_{a},\sigma_{a})-\Phi_{d}(t;\mu_{d},\sigma_{d}).$ (4)
We name such a basic model the Truncated Normal (TN) model. The TN model has parameters $\theta=(\mu_{a},\mu_{d},\sigma_{a},\sigma_{d})$, where $\mu_{a}$ and $\sigma_{a}$ denote the location time and the scale of the arrivals, respectively, and $\mu_{d}$ and $\sigma_{d}$, which denote the average time and the scale of the departures, respectively.
The top panel of Figure 5 shows an example with a density function for arrivals (in blue) and departures (in red). The bottom panel shows the corresponding CDFs (dashed lines) together with their difference (red solid line) as the aggregate result of cars parked in the parking lot.
Figure 5: Example of the TN model with data from the Vilanova car park. (top) probability density function (PDF) of car arrival (blue) and departure times (red), (bottom) corresponding CDFs (dashed lines) and car park occupancy curve (red think solid line).
The TN model provides a continuous temporal profile that can be easily adjusted if the available data consists of aggregated occupancies. Additionally, it provides a clear interpretability of the underlying mechanisms shaping the occupancy profile. Despite its simplicity, we will show that the TN model achieves predictive performance comparable to or superior to more complex approaches.
Figure 6: Example of TNL model with data for the Quatre Camins car park. The car park fills up completely between 8 and 8:30 (cyan dashed vertical lines), which generates an excess of cars (dash-dotted black lines) which would arrive at the station but have to park elsewhere.
4.2 Model with Occupancy Limit
We can extend the basic TN model to account for a capacity limit in the parking lot by incorporating a new parameter into the model. Consider the total number of cars arriving in one day and let $\tau$ be the fraction of those cars that fit in the parking lot. We define the Truncated Normal with Limit (TNL) model as
$f_{\text{TNL}}(t;\theta)$ $=\bar{\Phi}_{a}(t;\mu_{a},\sigma_{a},\tau)-\Phi_{d}(t;\mu_{d},\sigma_{d}),$
with $\text{with }~{}\bar{\Phi}_{a}(t;\mu_{a},\sigma_{a},\tau)=\min\left(\frac{\Phi_{a}(t;\mu_{a},\sigma_{a})}{\tau},1\right)$ (5)
where $\bar{\Phi}_{a}(\cdot)$ is a truncated and re-scaled version of $\Phi_{a}(\cdot)$ accounting for the fraction $\tau$ only. In other words $\bar{\Phi}_{a}(\cdot)$ uses only the part of $\Phi_{a}(\cdot)$ which is smaller than $\tau$.
The TNL model allows, not only to characterize the moment $t_{L}$ when the capacity limit is reached (i.e. $\Phi_{a}(t_{L};\mu_{a},\sigma_{a})=\tau$), but also to have an estimate of the number of cars which did not find a free parking space. This is very relevant in scenarios where informed decision-making is required to determine policies to resize crowded parking lots.
Figure 5 illustrates the TNL model for a day when the parking lot capacity limit was reached between 8 and 8:30 (vertical dashed line in cyan). The top panel shows an example of a density function for arrivals (in blue) and departures (in red). The arrival distribution is truncated when the parking fills up. The dashed-dotted black line of the arrival distribution corresponds to the estimated density of cars arriving afterwards which no longer fit in the parking lot. The bottom panel shows the corresponding modified CDFs $\bar{\Phi}_{a}$ (dashed line in blue) and $\Phi_{d}$ (dashed line in red) together with the difference $\bar{\Phi}_{a}-\Phi_{d}$ (red solid line), as the aggregate result of cars parked in the lot. Again, the proportion of surplus cars is indicated by the dashed-dotted line.
We will allow a different value for $\tau$ for every day $d$ of training data while using the same values for other parameters of the two distributions. This way the parameters of the arrival distribution are independent of the actual moment a parking lot reaches its capacity limit and can be used, not only to predict the moment when the capacity limit will be reached, but also to have an estimate of the number of cars which did not find a free parking space, as we will show later in Section 5.3.
4.3 Optimizing Model Parameters from Observed Data
In this section, we describe the procedure for estimating the model parameters using observed car park occupancy data.
4.3.1 Fitting the TN Model
For the basic TN model, the objective is to estimate the parameters $\boldsymbol{\theta}=(\mu_{a},\mu_{d},\sigma_{a},\sigma_{d})$ from a dataset consisting of $N$ daily occupancy profiles, $\mathcal{D}~{}=~{}\{o_{1:T}\}_{n=1}^{N}$, recorded at $T$ discrete timestamps [^2] In our case, $T=48$ since the dataset has a 30-minute time resolution.. To ensure consistency, all occupancy profiles are normalized such that $\sum_{i=1}^{T}o_{i}=1$.
If occupancy profiles exhibit stability within a group, meaning there is minimal variation across weekdays or weekends, we can reasonably assume that occupancy at time $t$ is independently and identically distributed. For simplicity, we model occupancies as Gaussian-distributed with mean $f_{\boldsymbol{\theta}}(t)$ and variance $\beta^{2}$. The likelihood of the parameters $\boldsymbol{\theta}$ is given by
$p(o_{t}^{(n)}|t,\boldsymbol{\theta},\beta)$ $=\frac{1}{\beta\sqrt{2\pi}}\exp\left(-\frac{1}{2\beta^{2}}\left(o_{t}^{(n)}-f_{\boldsymbol{\theta}}(t)\right)^{2}\right)$ (6)
$p(\mathcal{D}|\boldsymbol{\theta},\beta)$ $=\prod_{n=1}^{N}\prod_{t=1}^{T}p(o_{t}^{(n)}|t,\boldsymbol{\theta},\beta).$ (7)
The optimization can be carried by minimizing the sum of squares loss
$\mathcal{L}^{*}=\min\sum_{n=1}^{N}\sum_{t=1}^{T}\left(o_{t}^{(n)}-f_{\boldsymbol{\theta}}(t)\right)^{2},$ (8)
and the variance is given by $\beta^{2}=\mathcal{L}^{*}/NT$.
4.3.2 Fitting the TNL Model
For the TNL model with occupancy limit, the set of fitted parameters $\boldsymbol{\theta}$ also includes a set of threshold parameters $\{\tau_{i}\}_{i=1}^{N}$. Thus, we allow $\tau$ to take a different value for each training day $i$ while keeping the other parameters fixed. This approach ensures that the parameters of the arrival distribution remain independent of the exact moment a parking lot reaches capacity. As a result, the model can not only predict when the capacity limit will be reached but also estimate the number of cars that were unable to find a free parking space. Note, however, that this approach requires a different normalization of the data. In this case, all occupancy profiles are scaled such that $\max(o_{1:T})=1$. This normalization is essential for identifying the moments when the parking lots reach full capacity.
4.3.3 Test-Training Split
To train the model, we split the available data into training and test sets after the data-cleaning process. The test set is fixed at three weeks, while the remaining data is used for training. For most stations, this results in an approximate $30/70$ test-training split.
4.4 Alternative Approaches
In this section, we briefly describe two alternative methods against which we compare our proposed modelling approaches.
4.4.1 Average Activity Profile
A straightforward method for predicting patterns influenced by circadian rhythms is the use of average activity profiles, as employed in [11]. We adapt this approach by computing the average over normalized, aggregated data, distinguishing between three groups: weekdays (Monday to Thursday), Fridays, and weekends (Saturday and Sunday).
4.4.2 Linear Regression
As another baseline, we use a linear regression model to predict car park occupancy based on prior activity. Formally, given a dataset of $N$ normalized daily occupancy profiles, $\mathcal{D}=\{o_{1:T}\}_{n=1}^{N}$, we aim to predict $o_{y}$ using past observations $o_{1:x}$, where $x<y$. Since occupancy profiles represent cumulative arrivals and departures, we apply the diff function to obtain changes in occupancy at each 30-minute interval. Thus, we define the transformed input as $\hat{o}_{1:x}=\{0,\mathtt{diff}(o_{1:x})\}$.
The regression model optimizes the following objective:
$\mathcal{L}^{*}_{\text{lr}}=\min\sum_{n=1}^{N}\left(\sum_{t=1}^{x}\beta_{t}\hat{o}_{t}^{(n)}+\beta_{0}-o_{y}^{(n)}\right)^{2}.$ (9)
5 Results and Discussion
In this section, we first evaluate how well the model fits the data, followed by an analysis of its effectiveness in predicting or nowcasting parking lot occupancy at a given moment.
5.1 Model Fit Quality
We begin by assessing how well the model represents both the training and test data. For this, we use the mean relative proportional error, computed relative to the maximum occupancy of a parking lot. For clarity, we group the test data into three categories based on observed activity patterns (as illustrated in Figure 3): weekdays (Monday to Thursday), Fridays, and weekends (Saturday and Sunday).
5.1.1 Example Fits on Training Data
Figure 7: Example of the TN model fitted to training data for the Vilanova car park. The corresponding parameters of the arrival and departure TN distributions are $\mu_{a}=~{}$06:55h, $\mu_{d}=~{}$18:40 for weekdays, $\mu_{a}=~{}$07:02h, $\mu_{d}=~{}$17:26h for Fridays and $\mu_{a}=~{}$07:49h weekends. Note that for the TN model the data is normalized by the area under the curve.
Figure 8: Example of the TNL model fitted to training data for the Quatre Camins car park. The corresponding parameters of the arrival and departure distributions are $\mu_{arr}=~{}$07:31h, $\mu_{dep}=~{}$19:24h for weekdays, $\mu_{arr}=~{}$07:43h, $\mu_{dep}=~{}$18:29h for Fridays and $\mu_{arr}=~{}$10:10h and for weekends. Dashed lines indicate stdv of the model curve. The car park usually reaches its capacity limit from Monday to Friday between 8 and 8:30. Note that for the TNL model the data is normalized by the max value.
We first illustrate model fits on the training data using two example car parks. Figure 7 shows the extracted TN model curves for the Vilanova car park, while Figure 8 shows the TNL model curves, incorporating an occupancy limit, for the Quatre Camins car park. The latter reaches maximum capacity on most weekdays and Fridays (since for the weekend profile this does not happen the basic TN model is used there).
The dark grey lines represent actual occupancy profiles from the training data. The best-fitting model curves are shown as solid red lines, while blue dashed lines indicate the standard deviation of the model fit. All curves are normalized: in the TN model, by the area under the curve, and in the TNL model, by the maximum value. The latter normalization is necessary to estimate excess demand in a given car park. For Figure 8, the model curve is plotted using the average excess across all training curves. We observe the best fits for weekday profiles, while Fridays, especially weekends, exhibit greater variability.
The fitted TN model curves and their corresponding parameters are shown in Figure 9. The figure presents model curves for weekdays (left panel) and Fridays (right panel) across all eight car parks, ordered by the difference between the location parameters $\mu$ of the arrival and departure distributions (represented as circle markers). The interval corresponding to ±$\pm 1$ standard deviation $\sigma$ around $\mu$ is indicated by dashed lines and horizontal markers for both the arrival and departure distributions.
A smaller time difference between arrivals and departures is observed on Fridays, primarily due to earlier departure times. This pattern aligns with the common practice of shorter work schedules on Fridays, often without lunch breaks. The most pronounced example is the Prat del Llobregat car park, where the average departure time shifts from approximately 17:30 to 14:30—a three-hour advancement. We exclude the corresponding curves for weekends, which are characterized by more extreme parameter values and greater variability.
Figure 9: Model curves, with parameters for the TN model, for weekends (left panel) and Fridays (right panel) for the different car parks. The dots indicated the mean $\mu$ and the dashed line the corresponding standard deviation interval $[\mu-\sigma,\mu+\sigma]$ of the underlying TN distributions for arrival (right) and departure (left).
5.1.2 Evaluating Model Performance on Test Data
We now analyze the model’s performance by evaluating errors in the test data. Examples for two individual stations are given in Appendix A.
Figure 10: Average proportional errors for different types of days of fitting the trained base model to test data for the 8 carparks with sufficient data. Shaded areas indicate the average standard deviation of the error measure. Stations with * use the model with occupancy limit (TNL) for weekdays and Fridays.
Taking the averages of these errors on the testing days and aggregating them into the three weekday groups, we obtain the errors depicted in Figure 10. The four rightmost stations reach their occupancy limit during weekdays (including Fridays), and for these, we depict the error of the TNL model for the weekday and Friday profiles. The shaded areas represent the average daily standard deviation of the prediction errors.
We find that, apart from the four stations that show significant variability on Fridays (Prat Del Llobregat, Cerdanyola, Mollet, and Sant Boi, the latter also on weekends), the average error is always below 4%. The high error on weekends for the Sant Boi car park is caused by party-goers visiting nearby nightlife attractions. This results in an additional high-occupancy phase during the early morning hours, which is not well captured by the current models.
In general, the models perform better on weekdays from Monday to Thursday, which exhibit more regular activity compared to Fridays. The error is nearly always lower for the weekend profile, but this is due to the lower activity on these days relative to the stations’ maximum occupancy.
5.2 Prediction
Next, we analyze the predictive quality of our model. More precisely, at a given moment in time, we aim to predict the evolution of occupancy in the subsequent hours of the day. Again, we measure prediction quality using the relative absolute error with respect to the maximum occupancy of a car park and aggregate weekdays into three groups. We also compare the results with average profiles, which represent the typical activity in a given car park for the corresponding types of days in the training data.
We make predictions by adjusting the model and the average profile to match the observed occupancy at the initial time. To achieve this, we simply shift and rescale the corresponding curves to best fit the known data. Formally, given an occupancy profile $o_{1:T}$ at a certain hour of the day $h<T$, we determine the parameters $\beta_{0}$ and $\beta_{1}$ that minimize
$\mathcal{L}^{*}=\min\sum_{t=1}^{h}\left(o_{t}-\beta_{0}-\beta_{1}f(t)\right)^{2},$ (10)
where $f(t)$ represents either a model or the average activity profile, and $\beta_{0}$ corresponds to the number of cars parked at the station at the beginning of the day.
In the case of the TNL model with an occupancy limit, only the arrival distribution $\Phi_{a}$ is used for fitting. This allows us to calculate the excess number of cars, resulting in the following procedure:
$\mathcal{L}^{*}=\min\sum_{t=1}^{\min(h,t_{m})}\left(o_{t}-\beta_{0}-\beta_{1}\Phi_{a}(t;\mu_{a},\sigma_{a})\right)^{2},$ (11)
where $t_{m}$ is the time slot at which $o_{t}$ first reaches its peak value:
$t_{m}=\min\{t\mid o_{t}=\max(o_{1:h})\}$
In other words, we only use the phase up until the car park reaches its maximum capacity for fitting.
To determine the fraction $\tau_{i}$ of the arrival distribution that fits within the car park, we solve:
$\tau_{i}=\max\{\Phi_{a}(t;\mu_{a},\sigma_{a})\quad|\quad\beta_{0}+\beta_{1}\Phi_{a}(t;\mu_{a},\sigma_{a},\tau_{i})\leq\max(o_{1:T})\}.$ (12)
This can be simplified to:
$\Phi_{a}(t;\mu_{a},\sigma_{a})\leq\frac{\max(o_{1:T})-\beta_{0}}{\beta_{1}}$ (13)
which leads to:
$\tau_{i}=\frac{\max(o_{1:T})-\beta_{0}}{\beta_{1}}.$ (14)
The departure distribution $\Phi_{d}(t;\mu_{d},\sigma_{d})$ is then rescaled as follows:
$\bar{\Phi}_{d}(t;\mu_{d},\sigma_{d})=\Phi_{d}(t;\mu_{d},\sigma_{d})\cdot(\max(o_{1:T})-\beta_{0}).$ (15)
Essentially, this ensures that the departure distribution returns to a baseline of $\beta_{0}$ cars at the end of the day.
The excess, i.e., the number of cars that the arrival distribution predicts will arrive at the parking lot after it reaches full capacity, can be calculated as
$\mathrm{Excess}(T)=\beta_{1}\cdot\max(\Phi_{a}(t;\mu_{a},\sigma_{a}))+\beta_{0}-\max(o_{1:T})=\beta_{1}+\beta_{0}-\max(o_{1:T}).$ (16)
Figures 11 and 12 provide examples of predictions performed at 7:00, 15:00, and 19:00 for the Vilanova (TN model) and Quatre Camins (TNL model) car parks.
The left panels display real data (dash-dotted lines) along with the predicted occupancy, while the right panels show the corresponding prediction errors for both the model curve (in blue) and the average profile (in orange). The average errors are depicted as dashed lines. The dark gray area represents the data used to generate the prediction.
Figure 12 also includes an estimate of the surplus of cars that could not fit in the car park, indicated by the black continuous line. We observe that the performance of the TNL model is slightly better than that of the average profile for the Quatre Camins car park, and the surplus prediction remains stable across the three depicted time points.
Figure 11: Three examples for predictions in a car park (Vilanova) with the TN model. Left: The grey areas indicate the known activity used to predict the number of parked cars (red dashed-dotted lines) during the remaining hours of the day. Right: evolution of the relative prediction error in %. Blue lines indicate performance of the TN model while orange lines show the comparison with an average day-cycle profile.
Figure 12: Three examples for predictions in a car park which fills up (QuatreCamins) with the TNL model. Left: The grey areas indicate the known activity used to predict the number of parked cars (red dashed-dotted lines) during the remaining hours of the day. Right: evolution of the relative prediction error in %. Blue lines indicate performance of the TNL model while orange lines show the comparison with an average day-cycle profile. Black line gives an estimate of the surplus (how many cars will not or did not fit in the parking).
5.3 Nowcasting
To further evaluate the model’s performance, we apply it to a nowcasting task, i.e., we assess how well it predicts the number of cars in the car parks one hour ahead.
We quantify this by measuring the following prediction error, using the parameters obtained in Eq.(10):
$\mathcal{E}=\sum_{t=h}^{h+w}\frac{|o_{t}-\beta_{0}-\beta_{1}f(t)|}{w\cdot\max(o_{1:T})},$ (17)
where $w=2$ (corresponding to a prediction window of 1 hour with a 30-minute data resolution).
Figure 13 shows violin plots of the distribution of nowcasting errors, aggregated by car park and activity profile. The figure compares prediction errors between the basic model (darker colours) and the average activity profile (brighter colours).
Both approaches exhibit similar performance, with the average profile performing slightly better on average. As before, performance is worst on Fridays (orange tones), improves on weekdays from Monday to Thursday (blue tones), and is best on weekends (green tones), although the latter is likely caused by the lower number of parked cars on weekends in general.
Nowcasting is performed from 7:00 to 23:00 in 30-minute intervals, with predictions made one hour ahead.
Figure 13: Violin plots of the relative nowcasting error (%) for different parking lots, predicting the number of cars in the parking area during the next hour based on data available up to a given time. Darker violins represent the performance of the TN model, while brighter violins show comparisons with an average day-cycle profile. The brightest violin plots correspond to a simple linear regression. Nowcasting is performed from 7:00 to 23:00. Blue horizontal lines indicate the mean of the distributions, the white dot represents the median, and the black vertical bar spans the 25th to 75th percentiles. The legend includes a pairwise performance comparison with the TN model.
Figure 14: Violin plots of the relative nowcasting error (%) for different parking lots, predicting the number of cars in the parking area during the next hour based on data available up to a given time. The darkest violins represent the performance of the TNL model with thresholds, slightly brighter violins correspond to the baseline TN model, the second brightest violins represent the average day-cycle profile, and the brightest violins correspond to the Linear Regression model. Nowcasting is performed from 7:00 to 15:00. Blue horizontal lines indicate the mean of the distributions, the white dot represents the median, and the black vertical bar spans the 25th to 75th percentiles. The legend includes a pairwise performance comparison with the TNL model.
Figure 14 compares the performance of the model with activity thresholds for the four stations that reach their capacity limit. In this case, we display nowcasting errors only for the period between 7:00 and 15:00 (with 30-minute intervals). After 15:00, car parks begin to empty, reducing the advantage of using the threshold-based model.
Violin plots with intermediate colour tones represent the models with thresholds, which consistently outperform the other models on weekdays across all three stations. However, the improvement on Fridays is less pronounced, with only a slight advantage observed for the Quatre Camins car park.
The figures also include a pairwise comparison of prediction improvement (shown in the legend), where we count how often the TNL model outperforms the other two models in each nowcasting instance. For example, on weekdays at the Quatre Camins station, the TNL model performs better than the baseline model in 81% of cases and surpasses the average profile model in 79% of the tested instances.
For better comparison, Table 1 presents the numerical values of the corresponding median prediction errors.
We observe that the Linear Regression Model generally exhibits the highest error values, particularly on Fridays, suggesting that it may not effectively capture key behavioural patterns. In contrast, the TNL Model demonstrates superior performance in predicting parking occupancy, especially on weekdays, where it consistently yields lower error values than the other models.
Notably, it performs exceptionally well for the Sant Boi and Mollet car parks, achieving median prediction errors as low as 0.18% and 0.08%, respectively, possibly due to more regular commuter patterns.
Table 1: Median nowcasting error in % for the four stations depicted in Figure 14 for weekdays and Fridays comparing the Truncated Normal with Limit (TNL), basic (TN), and Average Profile (Av. Prof.) model with a Linear Regression (L-Reg.). Nowcasting between 7:00 and 15:00 performed in 30-minute intervals.
Car Park weekdays Fridays
TNL TN Av. Prof. L-Reg. TNL TN Av. Prof. L-Reg.
Sant Sadurni 2.16 2.67 2.26 3.59 4.55 4.17 4.33 7.74
Sant Boi 0.18 2.29 1.82 0.31 3.69 3.19 2.44 6.60
Quatre Camins 0.08 1.88 1.84 1.26 1.62 3.60 2.04 3.14
Mollet 1.09 2.29 2.38 2.92 2.70 3.96 2.03 7.01
The Basic Model, in comparison, generally performs worse than the TNL Model, often showing significantly higher errors. For example, at Sant Boi, its weekday error (2.29) is more than ten times higher than that of the TNL Model (0.18). This highlights the importance of considering capacity limits in the modelling approach for improved accuracy.
Errors tend to increase across all models on Fridays, suggesting greater variability in parking behaviour. Additionally, since the difference between the TNL and Basic Model becomes less pronounced, this may indicate that, due to lower demand, the capacity limit of car parks is less frequently reached on Fridays in most stations.
6 Conclusion
We have introduced a simple yet effective model for predicting car park occupancy in park-and-ride facilities. The model leverages truncated normal distributions to represent car arrival and departure times. This approach enables efficient and accurate parking demand prediction and estimation using only a small set of parameters.
A key advantage of the model is that it relies solely on aggregate data, eliminating the need to monitor individual arrivals and departures. Validation using data from Barcelona’s metropolitan area demonstrates the model’s robustness, even when applied to heterogeneous datasets. Its effectiveness is evident in both prediction and nowcasting tasks.
Furthermore, an extended version of the model, Truncated Normal with Limit (TNL), can estimate the additional parking spaces required to accommodate demand that remains unmet when car parks reach capacity. The extended model explicitly accounts for thresholding behaviour in station overflow, allowing it to determine the moment at which the capacity limit is reached and estimate the number of vehicles unable to find a free parking space.
This capability is particularly relevant for making informed decisions regarding the expansion of crowded parking lots. This estimation of unmet demand is a particular novelty of our study. To our knowledge, previous research has not directly addressed this issue, despite the fact that parking shortages rank among the top three reasons for not using Park-and-Ride facilities [14].
Additionally, our work differentiates itself from related studies through its specific focus on the modelling approach. Unlike some studies that prioritize prediction using techniques without an underlying behavioural model [3, 4, 5], our approach is more ambitious, aiming to understand the fundamental behavioural patterns of commuters rather than merely forecasting occupancy trends.
Moreover, the use of truncated normal distributions provides two key advantages over more complex, overparametrized models, such as neural networks [4]. First, it relies on a minimal number of parameters, ensuring computational efficiency. Second, these parameters correspond to behavioural metrics that are inherently interpretable, offering valuable insights into commuter patterns.
In contrast with the prevailing trend of requiring large-scale, fine-grained, high-precision datasets [2], our approach demonstrates that accurate predictions can be achieved using aggregated and limited data.
Ultimately, we believe this work will contribute to improved urban planning by offering a practical tool for understanding and predicting commuter behaviour in park-and-ride systems. By balancing simplicity, interpretability, and predictive accuracy, our approach can support data-driven decision-making for optimizing parking infrastructure and enhancing sustainable mobility solutions.
List of abbreviations
  • ARIMA: auto-regressive integrated moving average
  • ATM: Autoritat de Transport Metropolità
  • CDF: cumulative distribution function
  • PDF: probability density function
  • SVM: support vector machine
  • TN: Truncated Normal
  • TNL: Truncated Normal with Limit
Declarations
6.1 Availability of data and materials
All the data used in the paper and the corresponding code are publicly available at: https://github.com/aig-upf/car-park
6.2 Competing interests
The authors declare that they have no competing interests.
6.3 Funding
VG has received funding from “la Caixa” Foundation (ID 100010434), under the agreement LCF/PR/PR16/51110009.
6.4 Authors’ contribution
AK and VG designed the research project and performed data analysis, model optimization and evaluation. JF and DM performed data analysis, model optimization and evaluation. All authors contributed to the writing and reviewed the manuscript.
6.5 Acknowledgements
The authors would like to thank ATM Barcelona (Autoritat del Transport Metropolità) for providing the data.
Appendix A Error Analysis for Two Representative Car Parks
In this section, we present the relative errors of the two models for two example car parks.
For the TN model, we examine the relative errors depicted in Figure 15 for the Vilanova car park, aggregated by days of the week and hours of the day. The figure displays errors separately for each day of the week, with the average error per hour represented by red continuous lines and the corresponding standard deviations indicated by dashed blue lines. The black dashed-dotted line represents the average error across all hours of the day.
Similarly, Figure 16 provides an evaluation of the TNL model’s performance for the Quatre Camins car park. We observe how this model results in large periods with very small errors but exhibits larger errors in the afternoon during the departure phase of the cars, which is not influenced by the addition of a threshold to the model.
Figure 15: TN Model: Proportional errors for different weekdays when evaluating the trained model on test data for the Vilanova car park.
Figure 16: TNL Model: Proportional errors for different weekdays when evaluating the trained model with an occupancy limit on test data for the Quatre Camins car park. The mean individual training occupancy threshold was used as the model curve.